Find two consecutive positive integers such that the sum of their squares is 41.
Answer by algebrahouse.com(1659)

"Find two consecutive positive integers such that the sum of their squares is 41."

x = 1st positive integer
x + 1 = 2nd positive integer

x² + (x + 1)² = 41 {sum of squares is 41}
x² + (x + 1)(x + 1) = 41 {when you square a binomial, multiply it by itself}
x² + x² + 2x + 1 = 41 {used foil method}
2x² + 2x + 1 = 41 {combined like terms}
2x² + 2x - 40 = 0 {subtracted 41 from both sides}
2(x² + x - 20) = 0 {factored 2 out}
2(x + 5)(x - 4) = 0 {factored into two binomials}
x + 5 = 0 or x - 4 = 0 {set each factor equal to 0, excluding the 2}
x = -5 or x = 4 {solved each equation for x}
x = 4 {only positive integer}
x + 1 = 5 {substituted 4, in for x, into x + 1}

4 and 5 are the consecutive positive integers
www.algebrahouse.com